# TrialPanorama: Database and Benchmark for Systematic Review and Design of Clinical Trials

### v1

@@ 1 Introduction
Developing new drugs is a time-consuming and costly endeavor, often requiring up to a decade and exceeding \$2 billion in expenses [1]. A substantial portion of this burden stems from the human study phase, during which clinical trials are conducted to assess safety and efficacy across phases I through IV. In addition, clinicians, researchers, and trial designers routinely search and review clinical trial records and results. These efforts support critical objectives such as designing new trials, evaluating treatments, and informing clinical decision-making in evidence-based medicine [2, 3].
Clinical trial records are dispersed across global registries, including ClinicalTrials.gov [4], the International Clinical Trials Registry Platform (ICTRP) [5], the EU Clinical Trials Register [6], and others [7, 8]. Additionally, many trials are reported in the literature, particularly on PubMed. Prior research has focused on curating subsets of this data for specific objectives, such as clinical trial results [9, 10], drug and condition annotations [11, 12], trial-publication linkages [13], and structured outcomes [14]. A unified view that integrates these diverse aspects of clinical trial data can substantially improve evidence synthesis in systematic reviews and enable more informed, data-driven approaches to designing future clinical trials [15, 16, 17]. In this work, we build TrialPanorama, a clinical trial database that integrates multiple data sources and standardizes trial-level design and result attributes, aiming to provide a great resource for clinical trial design and review as well as developing AI models for these tasks.
Artificial intelligence (AI), particularly with the emergence of large language models (LLMs), has garnered significant attention in clinical trial applications. Recent advances have explored tasks including predicting clinical trial outcomes [18, 19, 20], matching patients to eligible trials [21, 22], extracting cohort real-world data [23, 24], writing clinical trial documents [25, 26], and accelerating systematic review of the literature of clinical studies [27, 28]. However, these developments are often limited by the lack of standardized, comprehensive clinical trial data suitable for AI development. TrialPanorama bridges this gap by (1) offering curated benchmark data to support the development and evaluation of AI methods for trial applications, and (2) providing source-grounded evidence to enhance AI for trial applications through retrieval-augmented generation (RAG) [29].
Table 1: Comparison of TrialPanorama and prior clinical trial datasets and benchmarks. Columns: Size = dataset size (# trials), Src = data source (CTG = ClinicalTrials.gov), Des = structured trial design data (e.g., arms, eligibility), Res = structured trial results (e.g., adverse events), Out = trial outcomes, Rel = trial-study/publication/review relations, Ont = linked to medical ontologies (e.g., DrugBank, MedDRA), Bench-D = trial design benchmark tasks, Bench-SLR = systematic literature review benchmark tasks.
Dataset / Benchmark Size Src Des Res Out Rel Ont Bench-D Bench-SLR
TrialPanorama (Ours) 1.6M Multi(15 sources)
TOP [19] 10K CTG
CTKG [11] 8K CTG
CTGResults [9] 56K CTG
CTO [14] 125K CTG
TrialBench [30] 400K CTG
CARE [10] 700 pubs PubMed
PubMedKG [31] 480K PubMed,CTG
TrialReviewBench [28] 100 reviews PubMed,CTG
LEADSInstruct [27] 21K reviews PubMed,CTG
In this work, we introduce TrialPanorama, the largest-to-date structured clinical trial database, integrating over two million records from 16 global sources, including clinical trial registries, biomedical literature, and systematic reviews. As illustrated in Figure 1a, the database captures essential elements of clinical trials, such as study designs, interventions, conditions, outcomes, and biomarkers, and standardizes them through mappings to biomedical ontologies like DrugBank and MedDRA.
Building on this data foundation, we further demonstrate how TrialPanorama enables tasks that support AI development for trial design and systematic review. These tasks span both upstream evidence synthesis (e.g., study search, screening, summarization) and downstream trial planning (e.g., arm design, eligibility criteria, endpoint selection, sample size estimation, and trial completion prediction), as illustrated in Figures 1b and  1c. Compared to existing trial datasets and benchmarks (Table 1), TrialPanorama offers broader coverage, richer structure, and stronger alignment with real-world AI use cases in clinical research. We release the database and benchmark to facilitate future research in this space.
Figure 1: Overview of the TrialPanorama dataset and benchmark suite. (a) Schema of the structured clinical trial database, showing linked tables for trial protocols, interventions, outcomes, and metadata. (b) Summary of the systematic review benchmark tasks. (c) Summary of the clinical trial design benchmark tasks.
2 TrialPanorama: Database
As illustrated in Figure 1a, the TrialPanorama database comprises ten main tables: studies, conditions, drugs, endpoints, relations, biomarkers, outcomes, adverse events, results, and relations. These tables are organized into four conceptual clusters: trial attributes, trial protocols, trial results, and study-level links. The trial metadata cluster captures core metadata such as study title, brief summary, sponsor, start year, and recruitment status. The trial protocol cluster describes the setup and design of the study, including tested drugs, targeted conditions, and patient group allocations. The trial results cluster encompasses reported findings, including the trial outcomes, adverse events, and efficacy results. The study-level link cluster encodes relationships across studies, such as mappings between registry records and corresponding publications, or reviews aggregating multiple studies on a common clinical topic. In the following subsections, we describe the content of each table in detail. Data acquisition and processing procedures are provided in Appendix A.
2.1 Trial metadata
The studies table serves as the central registry in TrialPanorama, containing 1,657,476 clinical trial records from 15 diverse sources, including ClinicalTrials.gov, PubMed, CHICTR, etc. The distribution of the sources is illustrated in Figure 2a. Each row, uniquely indexed by a study_id, captures metadata such as trial type (INTERVENTIONAL, OBSERVATIONAL), sponsor (e.g., university, hospital, pharmaceutical), and temporal attributes including start_year, enabling longitudinal analysis. The table also includes regulatory fields like phase (e.g., PHASE1PHASE4) and recruitment status (e.g., RECRUITING, COMPLETED, SUSPENDED), providing a standardized overview of the trial’s attributes.
2.2 Trial protocols
The trial protocol cluster in TrialPanorama captures structured information describing the design and setup of clinical studies, including target populations, interventions, outcomes, and biological markers. This cluster comprises five key tables: drugs, conditions, biomarkers, dispositions, and endpoints.
Figure 2: Overview of the database statistics. a, Distribution of clinical trials across major data sources. b, Frequency distribution of the top 100 drugs mentioned in the database, with the most common drug classes annotated. c, Frequency distribution of the top 100 medical conditions, annotated by condition categories.
Drugs.
The drugs table records the therapeutic interventions evaluated across trials, comprising 1,498,985 rows and 15 columns. Each entry is linked to a study_id, with over 616,000 unique studies represented. The distribution of top frequent drugs is shown in Figure 2b. The drug_name field captures 404,175 unique names, including both generic labels (e.g., “Placebo”) and brand/composite entries. To enable structured pharmacological analysis, drugs are linked to DrugBank identifiers (drugbank_id) and normalized RxNorm terms (rx_normalized_name). Mechanism of action (MoA) information is encoded via the drug_moa_id field, which maps each drug to its biological targets, such as enzymes, receptors, or transporters, and their interaction types (e.g., inhibitor, agonist). The definitions and details of each MoA are stored in a separate drug_moa table, where drug_moa_id serves as the key linking back to the drugs table.
Conditions.
The conditions table catalogs the disease or physiological focus of each study, containing 10,133,060 entries from over 1,074,458 studies. Figure 2c shows the most frequent conditions. The condition_name field includes 5,236 unique MeSH-derived terms representing medical conditions, while condition_mesh_id ensures mapping to standardized biomedical ontologies. These annotations facilitate the retrieval and grouping of studies by clinical focus (e.g., cardiovascular or neurological diseases).
Biomarkers.
The biomarkers table records therapeutic biomolecular and clinical markers used to stratify, monitor, or predict treatment response. It contains 477,717 rows across 339,894 unique studies. Each biomarker is annotated by its biomarker_type (e.g., predictive, surrogate endpoint, safety), biomarker_name, and associated gene markers (biomarker_genes). This table supports downstream applications such as biomarker-driven subgroup analysis or treatment personalization.
Dispositions.
The dispositions table describes how participants are grouped and assigned to interventions within each study. With over 4.5 million entries covering 1.47 million unique studies, it specifies the intervention_type (e.g., DRUG, DEVICE, BEHAVIORAL), group_type (e.g., EXPERIMENTAL, COMPARATOR, PLACEBO), and intervention_name. This table is critical for reconstructing arm-level trial structure and modeling allocation strategies across diverse study designs.
Endpoints.
The endpoints table defines the primary clinical outcomes of interest across 2,621,226 rows and 1.38 million trials. Each entry includes a primary_endpoint label (e.g., Overall Survival, Quality of Life), categorized into high-level primary_endpoint_domain (e.g., mortality, physiological) and fine-grained primary_endpoint_subdomain (e.g., cardiovascular outcomes, emotional well-being).
2.3 Trial results
The trial results cluster captures the observed clinical outcomes reported at the end of each study, including both efficacy and safety measures. It comprises three main tables: outcomes, results, and adverse_events, each offering distinct perspectives on the impact of tested interventions.
Outcomes.
The outcomes table contains 532,525 entries, each uniquely mapped to a study. It records overall trial status (overall_status), such as completed, recruiting, or terminated, and provides a high-level assessment of outcome success through the outcome_type field (e.g., completed, positive outcome/primary endpoint(s) met, unknown, or terminated, other/unclear). These annotations facilitate downstream analyses such as success rate estimation and status tracking across therapeutic areas and time.
Results.
The results table comprises over 5.7 million entries covering more than 1.1 million studies. It records arm-level trial outcomes, including structured descriptions of study populations (population), tested interventions (interventions), observed clinical outcomes (outcomes), and treatment group type (group_type). The granularity of eligibility criteria and baseline characteristics captured in the population field supports fine-grained subgroup analysis, while the group_type distinction (e.g., EXPERIMENTAL, COMPARATOR) enables comparative outcome interpretation across arms.
Adverse Events.
The adverse_events table documents reported safety events across 6,145,012 rows and over 530,000 unique studies. It captures adverse_event_name (e.g., Headache, Nausea, Diarrhoea) along with corresponding MedDRA-coded identifiers (meddra_id), enabling standardized aggregation and harmonization of safety profiles across studies. These data support pharmacovigilance and risk-benefit assessments for individual therapies or therapeutic classes.
2.4 Study links
The relations table captures cross-study relationships that connect clinical trial records from different sources or define higher-order linkages such as reviews or citation networks. This table contains 940,582 relation instances and supports integration, traceability, and synthesis of trial-level evidence. Each row represents a directed edge between a head_id and a tail_id, with associated head_type and tail_type identifying the origin of the studies (e.g., PubMed, ClinicalTrials.gov, or regional registries). This structure allows the database to represent connections between registry entries and publications describing the same study, or group related records across different national databases. The semantic type of each link is encoded in the relation_type field, which includes labels such as linked_to (e.g., registry-publication pair), cite (e.g., PubMed-to-PubMed citation), include (e.g., a review including a trial), and exclude (e.g., a review explicitly excluding a study). These relationships enable downstream use cases such as study clustering, citation graph construction, and synthesis of trial networks for systematic reviews.
Figure 3: Manual data quality assessment across key database tables. The y-axis shows the proportion of valid (i.e., correctly extracted) records. Sample sizes (n) are shown below each bar.
2.5 Data quality assessment
To assess the quality of TrialPanorama, we employed seven annotators to manually verify approximately 1,500 records sampled across all major tables, stratified by data sources. Each sampled record was evaluated against the original source to determine its validity. As summarized in Figure 3, the database exhibits high overall accuracy, with most tables showing a validity rate above 90%. Notably, errors observed in some tables, particularly conditions and biomarkers, can be attributed to challenges in text-based pattern matching and the use of LLMs for information extraction. Despite these limitations, the results highlight the feasibility and promise of using LLMs to automatically structure clinical trial data from heterogeneous and unstructured sources.
3 TrialPanorama: Benchmark
TrialPanorama serves as a valuable resource for a wide range of clinical trial applications by offering a global view of trial protocols, outcomes, and systematic reviews. Building on this comprehensive foundation, we develop a suite of representative benchmark tasks spanning two key application areas: three tasks focused on systematic evidence synthesis and five tasks supporting data-driven clinical trial design. These benchmarks are illustrated in Figures 1b and 1c.
3.1 Systematic review tasks
The PRISMA statement [32] outlines standard procedures for conducting systematic reviews and meta-analyses in medicine. Among these, three critical steps are: (1) study search: identifying candidate studies from literature databases; (2) study screening: selecting eligible studies based on inclusion and exclusion criteria; and (3) evidence synthesis: aggregating findings from included studies to answer a targeted clinical question, such as the efficacy of a therapy for a specific condition. Based on TrialPanorama, we construct benchmark tasks corresponding to each of these key steps: study search, study screening, and evidence summarization. By leveraging the relations table, we are able to identify linkages between systematic review articles and the clinical trials they include or exclude, enabling the construction of high-quality, labeled benchmarks for these tasks. We hold the most recent studies and reviews out as the test set. The details of benchmark building, setups, and example samples are in Appendix B.
Study Search.
The study search task evaluates an AI model’s ability to retrieve relevant clinical trials from a large corpus, given the structured context of a systematic review. It simulates the first PRISMA step: identifying candidate studies based on a review’s topic. Each benchmark instance includes a query review defined by its background, objectives, and selection criteria, and models must retrieve trials that were included in the final review. Ground-truth labels are derived from the relations table in TrialPanorama. The benchmark includes 8,004 training reviews linked to 175,577 studies and 100 test reviews with 4,324 relevant studies. Retrieval performance is measured using Recall@K.
Study Screening.
The study screening task evaluates an AI model’s ability to determine whether candidate studies meet the inclusion criteria of a systematic review, simulating the second PRISMA step of study selection. Each benchmark instance includes a structured review query (background, objectives, and criteria) and 10 candidate studies described by both unstructured text (title and abstract) and structured metadata from TrialPanorama (e.g., phase, intervention, conditions). Ground-truth labels are derived from explicit inclusion/exclusion links in the relations table. Notably, the excluded candidates represent hard negatives, as they were manually excluded by the original review authors and recorded in the review’s study list. The benchmark includes 7,931 training reviews with over 175,000 included and 229,000 excluded studies, and 100 test reviews with 491 included and 488 excluded studies. Performance is evaluated using accuracy, precision, and recall.
Evidence Summarization.
The evidence summarization task evaluates an AI model’s ability to answer clinical research questions by synthesizing findings from relevant studies, emulating the last PRISMA step: evidence summarization. Each instance includes (1) the review’s structured setup (background, objectives, selection criteria), (2) a set of study abstracts containing relevant or irrelevant evidence, and (3) a multiple-choice question summarizing possible conclusions. Input data and gold labels are constructed using the relations and results tables in TrialPanorama, where study-review links define the context and structured trial findings inform the correct answer. Answer choices reflect typical clinical conclusions (e.g., “limited evidence” or “effective with adverse effects”). The benchmark contains 4,636 reviews and 35,512 studies in the training set, and 1,000 reviews with 8,974 studies in the test set. Model performance is assessed using accuracy and macro-F1 score.
3.2 Clinical trial design tasks
Designing a clinical trial involves multiple interdependent components that collectively determine its scientific validity and operational feasibility. As illustrated in Figure 1c, key design tasks include: (1) sample size estimation, which determines the number of participants needed to ensure sufficient statistical power to detect treatment effects; (2) eligibility criteria design, which defines the characteristics of the target population eligible for inclusion; (3) arm design, which specifies the allocation of participants into treatment and control groups with defined interventions; and (4) endpoint design, which identifies the clinical measures used to assess efficacy and safety. In addition, trial completion assessment evaluates whether a trial is likely to reach completion or be terminated early due to design flaws, recruitment challenges, or safety concerns, thus providing a holistic perspective on the robustness of the trial design. We hold the most recent studies out as the test set. The details of benchmark building, setups, and example samples are in Appendix C.
Arm, Eligibility Criteria, and Endpoint Design.
The trial design tasks assess a model’s ability to identify optimal design components for a given trial protocol. Each task presents a structured input, consisting of the trial’s title and brief summary, and asks the model to select the correct choice from a set of candidate design elements, where distractors are sampled from similar trials to ensure difficulty. In the arm design task, the goal is to select the correct configuration of intervention groups; in eligibility criteria design, the correct inclusion and exclusion conditions; and in endpoint design, the appropriate outcome measures. Candidate options are drawn from structured records in TrialPanorama, including data from the studies, dispositions, conditions, and endpoints tables. For all three tasks, performance is evaluated using classification accuracy. The benchmark includes 27,321 training studies and 1,000 test studies for each of the three tasks.
Sample Size Estimation.
The sample size estimation task evaluates a model’s ability to infer the required number of participants for a clinical trial, based on its structured protocol, including target conditions, interventions, and primary outcomes, under a fixed drop-out assumption. Each instance draws data from the studies, dispositions, and endpoints tables in TrialPanorama, paired with gold-standard target values extracted from protocols or publications. The benchmark includes 1,143 training examples and 1,000 test studies. Model performance is assessed using two metrics: accuracy (percentage of estimation within ±20% of ground truth) and mean absolute error (MAE). This task requires both clinical reasoning and quantitative estimation skills grounded in trial design logic.
Trial Completion Assessment.
This task evaluates a model’s ability to predict whether a clinical trial will reach completion or terminate early based on its protocol, and, if terminated, to identify the most likely reason (e.g., efficacy, safety, enrollment failure). Each instance includes structured trial metadata, such as study phase, arm count, masking, allocation, and eligibility criteria, drawn from the studies, dispositions, and results tables in TrialPanorama. The benchmark comprises 42,195 training studies and 1,000 test studies. Model performance is evaluated by balanced accuracy on (1) the binary completion prediction and (2) multiclass termination reason classification. This task reflects real-world feasibility assessments during trial planning.
4 Experiment
We selected a diverse set of state-of-the-art proprietary and open-source LLMs for evaluation. These include GPT-4o and its lightweight variant GPT-4o-mini, as well as the recent reasoning-focused model O3-mini, all accessed via the Azure OpenAI Service. In addition, we include the LLaMA-3.3-70B-Instruct and LLaMA-3.1-8B-Instruct models, representing large and moderate-scale open-source LLMs, accessed through the Azure AI Inference Service. This selection allows us to compare models across different scales, architectures, and reasoning capabilities.
Table 2: Benchmark experiment results for systematic review tasks.
Tasks Study search Study screening Evidence summarization
Models Recall@100 Recall@500 Accuracy Precision Recall Accuracy Macro-F1
Llama-8B 11.6 16.0 55.3 46.8 58.2 70.4 71.1
Llama-70B 25.1 33.3 60.7 59.1 70.5 78.8 76.7
GPT-4o-mini 17.3 22.7 60.2 60.4 60.5 74.1 73.1
GPT-4o 22.9 27.3 66.1 64.4 72.3 72.8 65.2
O3-mini 27.6 37.8 71.6 71.9 71.3 72.7 53.8
4.1 Systematic review tasks
Table 2 presents the performance of various LLMs across three core systematic review tasks: study search, study screening, and evidence summarization. The results highlight both general limitations and comparative strengths across models. First, performance on the study search task is uniformly low, with Recall@100 ranging from 11.6 (LLaMA-8B) to 27.6 (O3-mini). This reflects a key weakness of general-purpose LLMs: they struggle to generate precise and discriminative search queries that can retrieve relevant trials from a large corpus. Notably, O3-mini outperforms both proprietary (GPT-4o, GPT-4o-mini) and open-source (LLaMA-70B, LLaMA-8B) models by a significant margin, suggesting its reasoning-focused training may better align with the query formulation challenge.
[…]

### v2

@@ Introduction
Developing new drugs remains a lengthy and expensive process that often spans close to a decade and costs more than two billion dollars [martin2017much]. A major share of this burden arises during the human study phase when clinical trials are conducted to evaluate safety and efficacy across phases I through IV. Beyond the execution of trials, clinicians, researchers, and trial designers routinely search for and review clinical trial protocols and results. These activities support essential objectives such as planning new studies, assessing treatments, and informing clinical decision making in evidence-based medicine [field2010meta, concato2017randomized].
There is growing interest in using artificial intelligence (AI), and especially large language models (LLMs), to improve the development and evaluation of clinical trials. Recent advances have explored a diverse set of tasks including predicting clinical trial outcomes [jin2020predicting, fu2022hint, wang2023spot], matching patients to appropriate trials [jin2024matching, wornow2025zero], extracting real world cohort data from eligibility criteria [yuan2019criteria2query, park2024criteria2query], drafting clinical trial documents [wang2023autotrial, wang2025informgen], and accelerating systematic review of clinical studies [wang2025foundation, wang2024accelerating]. Nonetheless, most depend on specialized data sources assembled from various specialized repositories of clinical trials. As a result, progress is constrained by the absence of standardized and comprehensive clinical trial resources suitable for AI development.
In this work, we introduce TrialPanorama, a clinical trial database and benchmark suite created from over one million records and fifteen global sources, to provide a broad foundation for trial design, trial review, and the development of AI for clinical trial tasks. We assemble clinical trial records from major global registries, including ClinicalTrials.gov [clinicaltrialsgov], the International Clinical Trials Registry Platform [whoictrp], the European Clinical Trials Register [euclinicaltrial], and several national registries [anzctr, chinesechictr]. We further incorporate trial-related evidence from literature sources such as PubMed. Prior studies have curated selected portions of these resources for focused objectives such as extracting trial results [shi2024constructing, naik2024care], annotating drugs and conditions [chen2022knowledge, brbic2024predicting], linking publications to registered trials [xu2020building], and generating structured outcome data [gao2024automatically]. Here, we put them into a unified representation that integrates these diverse elements, which can substantially strengthen evidence synthesis in systematic reviews and support data driven approaches to future trial planning [dunn2020time, holzhauer2023super, kim2017broadening].
Specifically, TrialPanorama is created as both a database and a tool suite for extracting clinical trial information and producing benchmark data for AI development in clinical trial applications. To illustrate its utility, we present a pipeline that extracts training and testing data from the resource and supports eight trial-related tasks. Five tasks focus on trial design and optimization, including arm design, eligibility criteria design, endpoint design, sample size estimation, and trial completion reasoning and prediction. Three tasks support systematic literature review workflows, including study search, study screening, and evidence question answering. We further evaluate open source and proprietary LLMs on these tasks to establish baseline performance and to demonstrate the value of TrialPanorama for future research.
Our benchmark further demonstrates the limitations of generic LLMs when applied to clinical trial tasks. We find that widely used LLMs display substantial gaps in study search, study screening, evidence summarization, and several trial design and optimization tasks. To explore the value of domain-specific data, we fine-tune two open-source models, Llama-3  [dubey2024llama] and Qwen-3  [yang2025qwen3], using training data extracted from TrialPanorama. These tuned models achieve competitive or superior performance relative to larger generic LLMs across both systematic review and trial design tasks. This improvement highlights the importance of high-quality trial-specific data foundations and demonstrates that TrialPanorama can meaningfully strengthen the capabilities of LLMs developed for clinical trial applications.
Results
TrialPanorama: a unified database and knowledge graph of clinical trials
Public information on clinical trials is available mainly through research publications and registry sites. For publications, we collected records from PubMed by focusing on papers labeled as randomized controlled trials and related tags, then applied a machine learning classifier to identify those that truly report clinical trial studies. For registry sources, we gathered data from major countries and regions worldwide, including the large registry clinicaltrials.gov and others such as CHICTR in China. Together, these sources yield more than 1.6 million trial records. As shown in Figure 1b, our final corpus includes 48.7% from PubMed, 31.6% from clinicaltrials.gov, and 19.6% from additional global registries. In addition, we collect the systematic literature review publications from PubMed and parse their references to obtain the PubMed ID of the cited clinical studies. Based on the sections of the references, we can further get the included and excluded studies from the review.
As shown in Figure 1a, for PubMed-sourced records, we gather abstracts and full texts when available, then use a large language model to extract core study attributes, including drugs, conditions, phase, primary endpoints, and related information. For global registry sources, we implement structured parsers to extract comparable fields so that all records can be transformed into a unified set of study attributes. Using this unified data, we apply a sequence of post-processing steps to standardize key fields. These include drugs, conditions, adverse events, biomarkers, and outcome measurements, which are normalized through entity extraction and matching with established databases such as DrugBank [drugbank], MedDRA [meddra], TheMarker [10.1093/nar/gkad862], RxNorm [RxNorm], MeSH term [meshterm], and COMET [kirkham2022core]. The entire pipeline integrates heuristic rules with LLM-based extraction.
As illustrated in Figure 1a, the goal of the post-processing stage is to construct a unified database of clinical trials. This includes a knowledge graph that links trials through shared target drugs, conditions, and other standardized entities, as well as connections based on shared systematic reviews, which often signal domain-relatedness. We also create a standardized relational database with multiple tables that capture detailed trial information, including studies, conditions, biomarkers, endpoints, adverse events, disposition, outcomes, and results. As shown in Table 1, we compare TrialPanorama against the most related databases from the literature and highlight its comprehensiveness in sources and the data field coverage. Figures 1c and 1d show the distributions of target drugs and conditions in TrialPanorama. We observe that a small set of agents appears with very high frequency, such as ethanol, ethinyloestradiol, paclitaxel, acetylsalicylic acid, and metformin, while the long tail consists of many agents used in only a modest number of trials. A similar pattern is present for conditions, where general categories related to pathologic processes, signs and symptoms, and broad disease classes appear most often, followed by a long tail of more specific disease groups. This unified resource enables efficient large-scale extraction and analysis of clinical trial information and serves as a strong data foundation for developing large language models that support clinical research tasks.
A pipeline for data extraction and LLM development
TrialPanorama enables efficient extraction of development data for large language models. Figure 2a illustrates an example workflow for constructing development data for the sample size estimation task. In Step 1, we identify the study fields that serve as inputs to the task. For this example, we use study-level attributes such as study type, year, phase, condition, and related descriptors as the trial setup, and use the field target accrual as the sample size value to be predicted. We further select the studies with full text so as to extract the statistical assumptions stated in the statistical analysis plan section. In Step 2, we organize these inputs into a standardized format for model development, which includes the instruction that specifies the task, the input that provides the trial setup as context, and the answer that contains the ground truth value for supervised learning. Through this approach, we can flexibly select study fields from the relational database and the knowledge graph to create instruction data tailored to many clinical research tasks.
As shown in Figures 2b and 2d, we use the TrialPanorama pipeline to extract development data for three systematic review tasks and five clinical trial design tasks. The systematic review tasks cover the core steps of a review workflow. The study search task takes a review topic as input and identifies the most relevant clinical trials from literature sources such as PubMed. The study screening task takes the review selection criteria and a candidate study set and evaluates the eligibility of each study for inclusion. The evidence summarization task takes the results of multiple clinical trials and produces a concise evidence summary that can support future clinical decision-making. For clinical trial design tasks, we include four tasks focused on study design fields: sample size estimation, criteria design, endpoint design, and arm design. We also include a trial completion assessment task that takes the trial setup as input, predicts whether the trial is likely to complete or terminate, and explains the factors that drive the prediction. Together, these tasks establish a comprehensive benchmark for large language models (LLMs) in clinical research settings and provide train-test splits for building vertical models tailored to clinical research applications. We hold out the latest studies as the test set to avoid the data leakage issues.
In Figure 2a, Step 3 applies a combination of supervised fine-tuning (SFT) and reinforcement learning with verifiable reward (RLVR) to adapt generic large language models to clinical research tasks. We first train models with SFT for all tasks, then apply RLVR to tasks that require deeper reasoning and offer a verifiable reward signal, such as sample size estimation. This staged approach enables broad task coverage while strengthening reasoning ability. Building on this framework, TrialPanorama benchmark provides a unified testbed to evaluate both open and proprietary models and to develop domain-adapted variants. For example, we adapt Qwen3-8B to clinical research by training it on the TrialPanorama corpus, producing Qwen3-8B-TP. As shown in Figure 2c, Qwen3-8B-TP delivers clear gains over the base Qwen3-8B across all tasks. Remarkably, this 8B vertical model surpasses Llama-70B, despite using about ten times fewer parameters, and also outperforms GPT-4o, underscoring the value of targeted post-training for clinical research workflows.
LLMs for clinical trial design and optimization
We present all results for trial design tasks in Figure 3. To benchmark LLM performance on these tasks, we include open source models Qwen3-8B [yang2025qwen3], Llama3-8B, and Llama3-70B [dubey2024llama], together with proprietary models GPT-4o-mini, GPT-4o, and the reasoning model o3-mini. For comparison, we also develop domain-adapted LLMs trained on TrialPanorama using the Qwen3-8B and Llama3-8B bases, which yield Qwen3-8B-TP and Llama3-8B-TP, where “TP” is short for TrialPanorama. Figures 3a and 3b illustrate the construction of the three trial design tasks. Each task presents the target trial setup on the left, including the disease area, treatment modalities, and study objective, followed by a multiple-choice question that mirrors common decisions made during protocol development. For arm design, models must determine which treatment arms best match the target trial, considering both the number of arms and the treatment composition. For eligibility criteria design, models must choose the inclusion and exclusion criteria that are most appropriate for the stated objective and patient population. For endpoint design, models must select the endpoint option that aligns with the clinical goal of the trial.
Figures 3c show the results of the three trial design tasks. Across arm design, eligibility criteria design, and endpoint design, proprietary frontier models outperform the open source baselines, while the domain-adapted models achieve the strongest overall accuracy. For arm design, the best open source model reaches 85.0%, whereas GPT-4o and O3-mini reach 86.3% and 89.0%, respectively. The domain-adapted models further improve performance, achieving 89.0% for Qwen3-8B-TP and 90.0% for Llama3-8B-TP. A similar pattern is observed in eligibility criteria design, where the baseline models plateau between 55.0% and 72.6%, while O3-mini reaches 84.4%. Domain adaptation again yields the highest accuracy, with Qwen3 8B TP at 87.1%. Endpoint design shows the largest gap between open source baselines and specialized models. Baseline models remain below 62.0%, whereas O3-mini reaches 69.1%. The domain-adapted variants achieve 74.3% and 78.2%, respectively, demonstrating consistent gains from tailoring the models to the clinical trial domain. These results indicate that domain-adapted LLMs provide the most reliable performance across diverse components of trial design.
Figure 3d shows the construction of the sample size estimation task, which requires multi-step statistical reasoning rather than pattern matching. The model must interpret a structured trial synopsis that includes the group definitions for control and experimental arms, along with the statistical assumptions needed for power analysis. Given this setup, the model must compute the required sample size that satisfies the primary power requirement. We assess model performance by computing the mean absolute error (MAE) of the model-estimated sample size against the original trial’s target accrual, which was made by human statisticians. As shown in Figure 3e, open source models exhibit large estimation errors, with mean absolute error ranging from 575.4 to 1769.6. Proprietary frontier models reduce the error substantially, reaching 411.4 for GPT-4o-mini and 452.1 for GPT-4o. Domain-adapted models achieve the best accuracy overall, with a mean absolute error of 356.3 for Llama3-8B-TP and 393.6 for Qwen3-8B-TP. These results highlight both the difficulty of this reasoning-heavy task and the effectiveness of targeted domain adaptation for improving statistical modeling performance.
Figure 3f shows the trial completion assessment task, where the model predicts whether a clinical trial will terminate early and, if so, selects the most plausible reason among options such as enrollment issues, safety concerns, lack of efficacy, or operational problems. The input includes structured trial attributes such as study type, phase, allocation, intervention model, eligibility criteria, outcome measurements, and arm design, requiring the model to integrate heterogeneous signals before making a binary prediction and generating a rationalization. As shown in the left panel, open source models achieve balanced accuracy between 52.2% and 56.4%, while GPT-4o reaches 59.3%. Domain-adapted models again deliver the highest performance, with Llama3-8B-TP reaching 74.3% and Qwen3-8B-TP reaching 72.8%. The right panel reports the rationalization accuracy, which is substantially more challenging. Baseline models remain below 26.0%, whereas the domain-adapted variants reach 26.0% for Llama3-8B-TP and 31.1% for Qwen3-8B-TP. These results indicate that domain adaptation improves both the termination prediction and the ability to identify the underlying reason, even though rationalization remains a challenging task.
LLMs for systematic literature review of clinical trials
Systematic literature review (SLR) is a foundational activity in clinical research, yet it remains one of the most labor-intensive and time-consuming stages of the trial lifecycle. In early trial planning, SLR supports evidence-grounded decisions by mapping the landscape of prior studies, identifying safety signals and adverse events in relevant patient groups, and establishing realistic benchmarks for study design. In later stages, SLR enables the integration of clinical evidence across historical trials to inform guideline development and clinical practice. To capture these essential workflows, we design a set of SLR tasks centered on three key steps: study search, study screening, and evidence summarization. These tasks, illustrated in Figure 4, provide a structured benchmark for evaluating how well LLMs can assist researchers throughout the clinical evidence synthesis pipeline.
Figure 4a illustrates the setup of the study search task. Given a review topic specified through the PICO framework, including the target population, intervention, comparison group, and outcome measures, the model must generate an effective search query for a literature database such as PubMed in order to retrieve as many relevant clinical studies as possible. We evaluate retrieval performance using Recall@100 and Recall@500, which measure the coverage of ground truth studies (i.e., the ones that were included in the review paper we reference in the test case) within the top retrieved results. As shown in Figure 4b, proprietary models and domain-adapted models achieve markedly stronger coverage than general open source baselines. GPT-4o and GPT-4o mini outperform Qwen3 and Llama models at both retrieval depths, and the specialized models trained on our dataset Qwen3-8B-TP and Llama3-8B-TP provide further gains. Under Recall@100, Qwen3-8B-TP reaches approximately 32% and Llama3-8B-TP reaches approximately 31%, compared with approximately 17% for Qwen3-8B and approximately 19% for Llama3-8B. Under Recall@500, Qwen3-8B-TP attains approximately 57%, and Llama3-8B-TP attains approximately 54%, while the strongest proprietary baseline GPT-4o reaches approximately 47%. These results show that targeted post-training on our clinical trial corpus substantially enhances an LLM’s ability to generate precise and comprehensive search queries for SLR workflows.
Figure 4c presents the setup of the study screening task. Given a set of papers retrieved for a review topic and the associated inclusion and exclusion criteria specified by the review, the model must decide for each paper whether it should be included or excluded. This task reflects the core filtering stage of an SLR workflow, where researchers determine which studies meet the methodological and clinical requirements of the review. We evaluate model performance using precision and recall, which together capture how accurately a model identifies eligible studies without introducing excessive false positives or false negatives. As shown in Figure 4d, proprietary models again surpass general-purpose open source LLMs, and the domain-adapted models yield the strongest overall performance. O3-mini and GPT-4o show clear gains over Qwen3 and Llama baselines, but the TP variants provide the largest improvements. Llama3-8B-TP achieves a precision of approximately 72%, and Qwen3-8B-TP reaches approximately 69%, compared with approximately 55% to 60% for the base open source models. In terms of recall, Qwen3-8B-TP attains approximately 86%, and Llama3-8B-TP reaches approximately 72%, again exceeding both open source baselines and proprietary models. These results indicate that targeted post-training substantially improves an LLM’s ability to apply complex inclusion and exclusion rules and reliably identify relevant studies for SLR workflows.
[…]